# TinyLLM: Learning a Small Student from Multiple Large Language Models

### v2

@@ Abstract
Transferring the reasoning capability from stronger large language models (LLMs) to smaller ones has been quite appealing, as smaller LLMs are more flexible to deploy with less expense. Among the existing solutions, knowledge distillation stands out due to its outstanding efficiency and generalization. However, existing methods suffer from several drawbacks, including limited knowledge diversity and the lack of rich contextual information. To solve the problems and facilitate the learning of compact language models, we propose TinyLLM, a new knowledge distillation paradigm to learn a small student LLM from multiple large teacher LLMs. In particular, we encourage the student LLM to not only generate the correct answers but also understand the rationales behind these answers. Given that different LLMs possess diverse reasoning skills, we guide the student model to assimilate knowledge from various teacher LLMs. We further introduce an in-context example generator and a teacher-forcing Chain-of-Thought strategy to ensure that the rationales are accurate and grounded in contextually appropriate scenarios. Extensive experiments on six datasets across two reasoning tasks demonstrate the superiority of our method. Results show that TinyLLM can outperform large teacher LLMs significantly, despite a considerably smaller model size.
1 Introduction
@@
Large language models (LLMs) have recently taken over various domains and web applications, including society  Rao et al. (2023), education Zelikman et al. (2023), and recommendations (Wu et al., 2023). Although cutting-edge language models like GPT-4 and Claude-2 have shown remarkable capability in producing coherent and contextually appropriate text, their smaller counterparts often fall short, especially in tasks that demand sophisticated reasoning and a deep level of understanding (Wei et al., 2022a). This discrepancy has been unveiled as the well-known scaling law of LLMs, which suggests a correlation between model size and reasoning, linguistic, and generalization capabilities (Kaplan et al., 2020). However, deploying these colossal models in a real-world setting poses significant challenges due to their computational requirements and resource demands, underscoring the importance of building efficient, smaller models that retain the power of their larger counterparts. Previous studies have shown that knowledge distillation is an instrumental tool in mitigating the performance gap between larger LLMs and smaller ones (Wan et al., 2023; Hsieh et al., 2023). Examples of effective distillation methods include DistilBERT Sanh et al. (2019), Alpaca Taori et al. (2023), and Vicuna (Zheng et al., 2023).
@@
However, existing methods suffer from two major drawbacks: (1) Limited Knowledge Diversity: Current research predominantly employs a single-teacher approach, which confines the learning scope of the student model to the knowledge derived from its own training and architecture designs (Ho et al., 2022; Magister et al., 2022; Li et al., 2023; Wang et al., 2022). This restricts the student model to a single perspective, potentially overlooking the diverse problem-solving strategies and reasoning capabilities exhibited by different models, limiting its breadth and depth of understanding. (2) Lack of Rich Contextual Information: While rationales play a vital role in effective reasoning Wei et al. (2022b); Kojima et al. (2022), current research primarily focuses on leveraging ground truth labels, which indicate the correct answer but do not provide insights into the reasoning and thought process behind that answer. In other words, learning the ground truth labels exclusively failed to capture the nuanced decision-making processes of the teachers, which are crucial for tasks requiring complex reasoning and interpretation.
@@
To solve these issues, we propose TinyLLM, a paradigm that facilitates the learning of a small student LLM by distilling knowledge from multiple large teacher LLMs with rationale guidance. Specifically, TinyLLM mitigates the limited knowledge diversity issue by involving multiple teacher models as co-advisors, which introduces a richer, varied knowledge source for the student to learn from. To fully exploit each teacher model and mitigate the lack of rich contextual information problem, TinyLLM asks the teacher for credible rationales to support the answers, thereby providing the student with a deeper understanding of the problem-solving process. By learning from multiple teachers, the student model can inherit a broader range of skills and knowledge, leading to better generalization capabilities. In addition, to ensure the rationales are grounded in contextually appropriate scenarios and reflect the true underlying reasoning procedure, TinyLLM features an in-context example generator and a teacher-forcing Chain-of-Thought strategy, making the teachers understand the task through demonstrations and therefore generate the accurate rationales.
@@
To thoroughly evaluate our approach, we conduct experiments on six datasets in commonsense and biomedical reasoning tasks. The results show that the usage of our paradigm enhances performance by +5.07% to +15.69% compared to full fine-tuning. Compared to the teacher models, TinyLLM achieve a superior performance improvement, e.g., up to +23.40% with significantly smaller model size, e.g., 1.1% to 26.0%. Furthermore, compared to the state-of-the-art distillation method, we improve the performance by +10.00% to +11.79% across different model sizes. In addition, we perform efficiency analyses, ablation studies, parameter sensitivities, and case studies to demonstrate and validate the effectiveness of the proposed method. To summarize, our main contributions are as follows:
@@
  • We identify two critical limitations in the existing knowledge distillation landscape for LLMs: 1) limited knowledge diversity and 2) lack of rich contextual information.
@@ 2 Related Work
Large Language Models. Recent advancements have seen the proposal of various Large Language Models (LLMs) Chung et al. (2022); Touvron et al. (2023a; b); Brown et al. (2020), which have showcased remarkable performance across a spectrum of tasks (Shi et al., 2023; Wei et al., 2024; Liu et al., 2024b; Tan et al., 2024; Liu et al., 2024a; Su et al., 2024). Central to these developments is question answering, a task that necessitates intricate reasoning and comprehensive understanding skills for text interpretation and generating suitable responses to queries (Lu et al., 2022; Zhu et al., 2021; Chen et al., 2023b; Tian et al., 2024). Despite their formidable learning capabilities, LLMs encounter limitations in accurately capturing factual knowledge and are prone to producing unsubstantiated responses (Zhao et al., 2023; Ji et al., 2023; Bang et al., 2023). Moreover, the extensive number of parameters within LLMs complicates their adaptation for downstream tasks (Workshop et al., 2022; Smith et al., 2022). To mitigate these challenges, several approaches aim to lessen the dependency on intensive training and reduce computational costs (Lester et al., 2021; Li & Liang, 2021; Hu et al., 2022). For example, Prompt Tuning Lester et al. (2021) employs soft prompts to adapt pre-trained LLMs for specific tasks.
@@
Chain of Thought. Recently, the use of rationales generated by LLMs has become a popular trend, setting itself apart from the traditional reliance on human-generated rationales Hase & Bansal (2022). Previously, human rationales have been used for model regularization Ross et al. (2017), as additional inputs for predictions Rajani et al. (2019), and to improve model performance Zaidan et al. (2007); Zhang et al. (2016); Camburu et al. (2018); Hancock et al. (2019); Pruthi et al. (2022). They also serve as gold standard labels for generating similar rationales to enhance interpretability Wiegreffe et al. (2021); Narang et al. (2020); Eisenstein et al. (2022). However, the cost of human rationales limits their widespread use. On the other hand, modern LLMs can generate high-quality reasoning steps to explain their predictions Wei et al. (2022b); Kojima et al. (2022), improving performance in few-shot or zero-shot learning Wei et al. (2022b); Kojima et al. (2022); Wang et al. (2022) and serving as self-improvement data (Zelikman et al., 2022; Huang et al., 2022). However, LLMs’ size hinders their deployment in practice. Correspondingly, recent research explores leveraging generated rationales for training smaller, task-specific models with minimal computational and memory overhead (Wang et al., 2021; Ho et al., 2022; Magister et al., 2022; Li et al., 2023). For example, PINTO Wang et al. (2022) presents an LLM pipeline that rationalizes via prompt-based learning. However, they still rely on an LLM for rationale generation at test-time, not fully addressing deployment challenges. In this work, we propose a multi-task learning paradigm with superior chain-of-thought reasoning capabilities, avoiding the dependence on teacher models during the test phase.
Knowledge Distillation. Recent LLMs such as PaLM 540b Chowdhery et al. (2023) present formidable challenges in terms of inference and fine-tuning, primarily attributable to the extensive computational resources they necessitate. This dependency and requirement on computation underscore the pivotal role of knowledge distillation Hinton et al. (2015); Tian et al. (2023b; a); Gou et al. (2021); Cho & Hariharan (2019), which has proved to alleviate the resource limitation by training a smaller model to mimic the large teacher model. In addition, the employment of the Chain-of-Thought paradigm has facilitated the generation of deliberative reasoning samples from the teacher models (Ho et al., 2022), allowing student models to concurrently grasp both the answers and the intricate reasoning of the teachers. This process strengthens the student model through multi-task learning endeavors (Hsieh et al., 2023). Correspondingly, efforts have been made to generate various rationales for each inquiry, seeking to ensure consistency in the predictions (Chen et al., 2023a). However, depending on the rationales from a single teacher model introduces bias and compromises thoroughness, thus not fully tapping into the capabilities of multi-teacher learning paradigms. These paradigms, by their very nature, hold the potential to enhance knowledge diversity, an aspect that remains largely unexplored.
3 Method
@@
Figure 1: Pipeline of TinyLLM: Given an input question, we first generate in-context examples and obtain rationales from multiple large LLMs via a teacher-forcing Chain-of-Thought strategy. Later, a small student LLM is trained to integrate rationales from different teachers via multi-task instruction tuning, along with the ground truth label.
3.1 Preliminary
@@
3.2 Obtaining Rationales from Teachers
In-context Example Generator. To enable the rationales that are generated by teachers to be grounded in contextually appropriate scenarios, we introduce an optional in-context example generator. This tool is designed to produce in-context examples for any given input, providing more detailed information about the input data and task. For simplicity, we select the examples randomly within the same dataset. This aids the teacher LLMs in comprehending the nature and specifics of the task more deeply. By integrating this generator, we facilitate a more informed and nuanced generation of rationales by the teacher models, enhancing the learning experience for the student model.
Teacher-forcing Chain-of-Thought. In addition, we design a teacher-forcing strategy to ensure the validity of the rationales. Compared to existing methods that simply employ a chain-of-thought (CoT) mechanism Wei et al. (2022b); Kojima et al. (2022) that an LLM is prompted with sets of questions and options $\{Q_{i},O_{i}\}$ to elicit rationales $R_{i}$ directly, TinyLLM posits a distinct advantage in integrating the correct answer $A_{i}$ into the input. We hypothesize that the inclusion of $A_{i}$ alongside $Q_{i}$ and $O_{i}$ facilitates a more nuanced understanding of the input context and the correct logical rationales leading to the answer, thereby facilitating a more informed and accurate generation process. Specifically, we consider the concatenation of questions, options, and answers $\{Q_{i},O_{i},A_{i}\}$ as the input to LLMs, enabling the generation of credible rationales.
@@ 3.2 Obtaining Rationales from Teachers
$R_{i}^{m}=T^{m}(Q_{i},O_{i},A_{i},P_{i};\theta_{T^{m}}).$ (1)
3.3 Learning a Small Student
A straightforward strategy to incorporate rationales as supervision is to append each rationale $R_{i}^{m}$ generated by the teacher models as supplementary input to the student model, along with the question $Q_{i}$ and options $O_{i}$. However, this method faces challenges due to limitations in computational resources at the inference stage, especially because rationales must be pre-generated for every data sample in both training and test sets (Wang et al., 2022). To overcome this issue, we employ rationales as a form of supervisory signal during the training process to develop a model that is adept at generating its explanations. Subsequently, this trained model can be utilized on the test set, eliminating the need for pre-generated rationales to facilitate accurate reasoning. Specifically, TinyLLM integrates rationales from multiple teacher models into a unified multi-task instruction tuning framework. This necessitates the assignment of a unique prefix $p$ for distinguishing between learning tasks from different teachers. The student model is trained not only to predict labels but also to generate rationales akin to those produced by the teachers. Accordingly, the overall loss function $\mathcal{L}$ is as follows:
$\mathcal{L}=\mathcal{L}_{A}+\sum_{m=1}^{M}\alpha^{m}\mathcal{L}_{T^{m}},$ (2)
@@ 3.3 Learning a Small Student
$\mathcal{L}_{A}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(Q_{i},O_{i},p_{A};\theta_{S}),A_{i}),$ (3)
$\mathcal{L}_{T^{m}}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(Q_{i},O_{i},p_{m};\theta_{S}),R_{i}^{m}),$ (4)
@@
4 Experiments
In this section, we rigorously test TinyLLM against a series of empirical benchmarks across varied datasets and reasoning tasks. In addition, we conduct efficiency analyses, ablation studies, parameter sensitivities, and case studies to demonstrate the effectiveness and superiority of our method.
4.1 Experimental Setup
Datasets. For the task of commonsense reasoning, we use OpenBookQA (OBQA) Mihaylov et al. (2018), The AI2 Reasoning Challenge (ARC) Clark et al. (2018), Physical Interaction Question Answering (PIQA) Bisk et al. (2020), and RiddleSense (Riddle) Lin et al. (2021). For the task of biomedical reasoning, we consider PubMedQA (PQA) Jin et al. (2019) and BioASQ (Tsatsaronis et al., 2015).
Baselines. We benchmark TinyLLM against the teacher’s performance and various baseline methods, including Inference-only that only leverages the pre-trained model for evaluation without training, and multiple fine-tuning methods that provide further adaptation. In particular, we consider LoRA Hu et al. (2022), full fine-tuning, and the PINTO method (Wang et al., 2022) for the fine-tuning methods.
We also compare TinyLLM with various knowledge distillation strategies. To illustrate, we include standard KD Hinton et al. (2015) that enforces the student to mimic the teacher’s labels and the Distill-step-by-step method (Hsieh et al., 2023) that leverages rationales.
Implementation Details. For all distillation baselines and TinyLLM, we set the learning rate to $5\times 10^{-5}$, batch size to 8, maximum input length to 1024, and epoch to 1. For Distill-step-by-step and TinyLLM, the trade-off weights $\alpha_{T_{n}}$ are explored within {0.01, 0.1, 0.5, 1, 2, 3}. We report the best result for Distill-step-by-step by leveraging different teacher models. For the choice of LLMs, we use FLAN-T5 Chung et al. (2022) small (80M), base (250M), and large (780M) as the student, and FLAN-T5 xlarge (3B) and LLaMA 2-chat Touvron et al. (2023b) (7B) as teachers. Experiments are conducted on four NVIDIA Tesla H100 GPUs.
4.2 Performance Comparison
@@
Method OBQA ARC PIQA Riddle PQA BioASQ Total
3B/7B Teacher
FLAN-T5 xlarge 69.20 68.24 58.43 53.73 71.50 65.85 64.49
@@
80M Student; Size: 2.7%/1.1%
Inference 16.60 19.31 20.78 13.33 38.00 47.97 26.00
OOD PINTO 46.40 26.87 48.10 25.29 60.00 80.49 47.86
@@
250M Student; Size: 8.3%/3.6%
Inference 31.00 23.00 30.47 30.78 48.00 57.72 36.83
OOD PINTO 50.40 38.63 52.12 34.90 61.75 82.93 53.46
@@
780M Student; Size: 26.0%/11.1%
Inference 50.40 51.07 51.90 39.80 64.25 63.41 53.47
OOD PINTO 62.20 52.10 57.13 42.94 70.00 84.55 61.49

### v3

# Beyond Answers: Transferring Reasoning Capabilities to Smaller LLMs Using Multi-Teacher Knowledge Distillation

@@ Abstract.
Transferring the reasoning capability from stronger large language models (LLMs) to smaller ones has been quite appealing, as smaller LLMs are more flexible to deploy with less expense. Among the existing solutions, knowledge distillation stands out due to its outstanding efficiency and generalization. However, existing methods suffer from several drawbacks, including limited knowledge diversity and the lack of rich contextual information. To solve the problems and facilitate the learning of compact language models, we propose TinyLLM, a new knowledge distillation paradigm to learn a small student LLM from multiple large teacher LLMs. In particular, we encourage the student LLM to not only generate the correct answers but also understand the rationales behind these answers. Given that different LLMs possess diverse reasoning skills, we guide the student model to assimilate knowledge from various teacher LLMs. We further introduce an in-context example generator and a teacher-forcing Chain-of-Thought strategy to ensure that the rationales are accurate and grounded in contextually appropriate scenarios. Extensive experiments on six datasets across two reasoning tasks demonstrate the superiority of our method. Results show that TinyLLM can outperform large teacher LLMs significantly, despite a considerably smaller model size. The source code is available at: https://github.com/YikunHan42/TinyLLM.
Knowledge Distillation, Knowledge Reasoning, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: 3701551.3703577conference: the 18th ACM International Conference on Web Search and Data Mining; March 10-14, 2025; Hannover, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Machine learning
1. Introduction
@@
Large language models (LLMs) have recently taken over various domains and web applications, including society (Rao et al., 2023), education (Zelikman et al., 2023), and recommendations (Wu et al., 2023). Although cutting-edge language models like GPT-4 and Claude-2 have shown remarkable capability in producing coherent and contextually appropriate text, their smaller counterparts often fall short, especially in tasks that demand sophisticated reasoning and a deep level of understanding (Wei et al., 2022a). This discrepancy has been unveiled as the well-known scaling law of LLMs, which suggests a correlation between model size and reasoning, linguistic, and generalization capabilities (Kaplan et al., 2020). However, deploying these colossal models in a real-world setting poses significant challenges due to their computational requirements and resource demands, underscoring the importance of building efficient, smaller models that retain the power of their larger counterparts. Previous studies have shown that knowledge distillation is an instrumental tool in mitigating the performance gap between larger LLMs and smaller ones (Wan et al., 2023; Hsieh et al., 2023). Examples of effective distillation methods include DistilBERT (Sanh et al., 2019), Alpaca (Taori et al., 2023) and Vicuna (Zheng et al., 2023).
@@
However, existing methods suffer from two major drawbacks: (1) Limited Knowledge Diversity: Current research predominantly employs a single-teacher approach, which confines the learning scope of the student model to the knowledge derived from its own training and architecture designs (Ho et al., 2022; Magister et al., 2022; Li et al., 2023; Wang et al., 2022a). This restricts the student model to a single perspective, potentially overlooking the diverse problem-solving strategies and reasoning capabilities exhibited by different models, limiting its breadth and depth of understanding. (2) Lack of Rich Contextual Information: While rationales play a vital role in effective reasoning (Wei et al., 2022b; Kojima et al., 2022), current research primarily focuses on leveraging ground truth labels, which indicate the correct answer but do not provide insights into the reasoning and thought process behind that answer. In other words, learning the ground truth labels exclusively failed to capture the nuanced decision-making processes of the teachers, which are crucial for tasks requiring complex reasoning and interpretation.
@@
To solve these issues, we propose TinyLLM, a paradigm that facilitates the learning of a small student LLM by distilling knowledge from multiple large teacher LLMs with rationale guidance. Specifically, TinyLLM mitigates the limited knowledge diversity issue by involving multiple teacher models as co-advisors, which introduces a richer, varied knowledge source for the student to learn from. To fully exploit each teacher model and mitigate the lack of rich contextual information problem, TinyLLM asks the teacher for the credible rationales to support the answers, thereby providing the student with a deeper understanding of the problem-solving process. By learning from multiple teachers, the student model can inherit a broader range of skills and knowledge, leading to better generalization capabilities. In addition, to ensure the rationales are grounded in contextually appropriate scenarios and reflect the true underlying reasoning procedure, TinyLLM features an in-context example generator and a teacher-forcing Chain-of-Thought strategy, making the teachers understand the task through demonstrations and therefore generate the accurate rationales.
@@
To thoroughly evaluate our approach, we conduct experiments on six datasets in commonsense and biomedical reasoning tasks. The results show that the usage of our paradigm enhances performance by +5.07% to +15.69% compared to full fine-tuning. Compared to the teacher models, TinyLLM achieves superior performance improvement, e.g., up to +23.40% with significantly smaller model size, e.g., 1.1% to 26.0%. In addition, compared to the state-of-the-art distillation methods, our approach improves the distillation performance by +10.00% to +11.79% across different model sizes. With the aim of further validating the effectiveness of our method, we perform efficiency analyses, ablation studies, parameter sensitivities, and case studies to provide a comprehensive evaluation across multiple dimensions. To summarize, our main contributions are as follows:
@@
  • We identify two critical limitations in the existing knowledge distillation landscape for LLMs: 1) limited knowledge diversity, and 2) lack of rich contextual information.
@@ 2. Related Work
Large Language Models. Recent advancements have seen the proposal of various Large Language Models (LLMs) (Raffel et al., 2020; Touvron et al., 2023a, b; Brown et al., 2020; Dubey et al., 2024), which have showcased remarkable performance across a spectrum of tasks (Shi et al., 2023; Wei et al., 2024; Liu et al., 2024b; Tan et al., 2024; Wang et al., 2023; Liu et al., 2024a). Central to these developments is question answering, a task that necessitates intricate reasoning and comprehensive understanding skills for text interpretation and generating suitable responses to queries (Lu et al., 2022; Zhu et al., 2021; Chen et al., 2024; Tian et al., 2024). Despite their formidable learning capabilities, LLMs encounter limitations in accurately capturing factual knowledge and are prone to producing unsubstantiated responses (Zhao et al., 2023; Ji et al., 2023; Bang et al., 2023). Moreover, the extensive number of parameters within LLMs complicates their adaptation for downstream tasks (Workshop et al., 2022; Smith et al., 2022). To mitigate these challenges, several approaches aim to lessen the dependency on intensive training and reduce computational costs (Lester et al., 2021; Li and Liang, 2021; Hu et al., 2022; Wang et al., 2024). For example, Prompt Tuning (Lester et al., 2021; Liu et al., 2021; Gu et al., 2021; Wang et al., 2022b) employs soft prompts to adapt pre-trained LLMs for specific tasks.
@@
Chain of Thought. Recently, the use of rationales generated by LLMs has become a popular trend, setting itself apart from the traditional reliance on human-generated rationales (Hase and Bansal, 2022; Huang and Chang, 2022). Previously, human rationales have been used for model regularization (Ross et al., 2017), as additional inputs for predictions (Rajani et al., 2019), and to improve model performance (Zaidan et al., 2007; Zhang et al., 2016; Camburu et al., 2018; Hancock et al., 2019; Pruthi et al., 2022; Fu et al., 2023; Huang et al., 2023; Zhang et al., [n. d.]). They also serve as gold standard labels for generating similar rationales to enhance interpretability (Wiegreffe et al., 2021; Narang et al., 2020; Eisenstein et al., 2022; Turpin et al., 2024). However, the cost of human rationales limits their widespread use. On the other hand, modern LLMs can generate high-quality reasoning steps to explain their predictions (Wei et al., 2022b; Kojima et al., 2022), improving performance in few-shot or zero-shot learning (Wei et al., 2022b; Kojima et al., 2022; Wang et al., 2022a) and serving as self-improvement data (Zelikman et al., 2022; Huang et al., 2022). However, LLMs’ size hinders their deployment in practice. Correspondingly, recent research explores leveraging generated rationales for training smaller, task-specific models with minimal computational and memory overhead (Wang et al., 2021; Ho et al., 2022; Magister et al., 2022; Li et al., 2023). For example, PINTO (Wang et al., 2022a) presents an LLM pipeline that rationalizes via prompt-based learning. However, they still rely on an LLM for rationale generation at test-time, not fully addressing deployment challenges. In this work, we propose a multi-task learning paradigm with superior chain-of-thought reasoning capabilities, avoiding the dependence on teacher models during the test phase.
Figure 1. Pipeline of TinyLLM: Given an input question, we first generate in-context examples and obtain rationales from multiple large LLMs via a teacher-forcing Chain-of-Thought strategy. Later, a small student LLM is trained to integrate rationales from different teachers via multi-task instruction tuning, along with the ground truth label.
Knowledge Distillation. Recent advancements in LLMs, such as PaLM 540B (Chowdhery et al., 2023; Anil et al., 2023) and LLaMA 3.1 405B (Dubey et al., 2024), bring formidable challenges in terms of inference and fine-tuning due to their immense computational requirements. Knowledge distillation (Hinton et al., 2015; Gou et al., 2021; Tian et al., 2023b, a; Cho and Hariharan, 2019; Guo et al., 2023; Agarwal et al., 2024; Xu et al., 2024) has emerged as a pivotal approach to mitigate these challenges by training smaller student models to mimic the behavior of larger teacher models, significantly reducing resource demands. White-box knowledge distillation leverages the output distributions and hidden states of teacher models to provide the student model with richer learning signals (Gu et al., 2024; Team et al., 2024; Ko et al., [n. d.]). However, this approach is infeasible for black-box settings where access to internal teacher model states is restricted. The Chain-of-Thought paradigm has further advanced knowledge distillation by enabling the generation of detailed reasoning samples from teacher models (Ho et al., 2022). These samples allow student models to learn both the correct answers and the intricate reasoning processes behind them. (Hsieh et al., 2023). To enhance consistency, recent efforts have focused on generating diverse rationales for the same query (Chen et al., 2023; Liu et al., 2023), ensuring more robust predictions. However, relying on rationales from a single teacher model risks introducing bias and limiting the scope of reasoning diversity. Multi-teacher learning paradigms, which integrate knowledge from multiple teacher models, hold the potential to overcome these limitations by enhancing knowledge diversity and capturing a broader range of reasoning styles.
3. Method
@@
3.1. Preliminary
@@
3.2. Obtaining Rationales from Teachers
To enforce the teacher LLMs generate high-quality rationales for student learning, our approach incorporates an In-context Example Generator to enrich contextual information, a Teacher-forcing Chain-of-Thought strategy to avoid misinformation, and rationale generation from multiple teachers to support student training.
In-context Example Generator. To enable that rationales generated by teacher models are grounded in contextually appropriate scenarios, we introduce an optional in-context example generator. This tool provides additional context by generating examples that include both questions and corresponding rationales in a zero-shot setting, enhancing the teacher models’ comprehension of task-specific nuances. Each generated in-context example offers a unique perspective on the dataset and provides additional insights. By including a range of in-context examples and sending them with the input question to the teacher LLMs, we enhance the input with richer information that aids the model in generating higher-quality rationales. This approach allows the student model to learn not only from correct answers but also from the underlying reasoning, thereby enhancing both the accuracy and interpretability of the distilled model.
Teacher-forcing Chain-of-Thought. In addition, we design a teacher-forcing strategy to ensure the validity of the rationales. Compared to existing methods that simply employ regular chain-of-thought (CoT) mechanisms (Wei et al., 2022b; Kojima et al., 2022), wherein an LLM is prompted with sets of questions and options $\{Q_{i},O_{i}\}$ to elicit rationales $R_{i}$ directly, TinyLLM posits a distinct advantage in integrating the correct answer $A_{i}$ into the input. We hypothesize that the inclusion of $A_{i}$ alongside $Q_{i}$ and $O_{i}$ facilitates a more nuanced understanding of the input context and the correct logical rationales leading to the answer, thereby facilitating a more informed and accurate generation process. Specifically, we consider the concatenation of questions, options, and answers $\{Q_{i},O_{i},A_{i}\}$ as the input to LLMs.
@@ 3.2. Obtaining Rationales from Teachers
(1) $R_{i}^{m}=T^{m}(Q_{i},O_{i},A_{i},P_{i};\theta_{T^{m}}).$
3.3. Learning a Small Student
A straightforward strategy to incorporate rationales as supervision is to append each rationale $R_{i}^{m}$ generated by the teacher models as supplementary input to the student model, along with the question $Q_{i}$ and options $O_{i}$. However, this method faces challenges due to limitations in computational resources at the inference stage, especially because rationales must be pre-generated for every data sample in both training and test sets (Wang et al., 2022a). To overcome this issue, we employ rationales as a form of supervisory signal during the training process to develop a model that is adept at generating its explanations. Subsequently, this trained model can be utilized on the test set, eliminating the need for pre-generated rationales to facilitate accurate reasoning. Specifically, TinyLLM integrates rationales from multiple teacher models into a unified multi-task instruction tuning framework. This necessitates the assignment of a unique prefix $p$ for distinguishing between learning tasks from different teachers. The student model is trained not only to predict labels but also to generate rationales akin to those produced by the teachers. Accordingly, the overall loss function $\mathcal{L}$ is as follows:
(2) $\mathcal{L}=\mathcal{L}_{A}+\sum_{m=1}^{M}\alpha^{m}\mathcal{L}_{T^{m}},$
@@ 3.3. Learning a Small Student
(3) $\mathcal{L}_{A}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(Q_{i},O_{i},p_{A};\theta_{S}),A_{i}),$
(4) $\mathcal{L}_{T^{m}}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(Q_{i},O_{i},p_{m};\theta_{S}),R_{i}^{m}),$
@@
Table 1. Overall experimental results. The best results across different datasets and LLM sizes are highlighted in bold. $\Delta_{FF}$ and $\Delta_{Distill}$ represent the relative performance improvement of TinyLLM to Full Fine-Tuning and Distill-step-by-step, respectively. Accuracy is used as the evaluation metric.
@@
Setting Method OBQA ARC PIQA Riddle PQA BioASQ Total
3B/7B Teacher FLAN-T5 xlarge 69.20 68.24 58.43 53.73 71.50 65.85 64.49
@@
80M Student Size: 2.7%/1.1% Inference 16.60 19.31 20.78 13.33 38.00 47.97 26.00
PINTO 46.40 26.87 48.10 25.29 60.00 80.49 47.86
@@
250M Student Size: 8.3%/3.6% Inference 31.00 23.00 30.47 30.78 48.00 57.72 36.83
PINTO 50.40 38.63 52.12 34.90 61.75 82.93 53.46
@@
780M Student Size: 26.0%/11.1% Inference 50.40 51.07 51.90 39.80 64.25 63.41 53.47
PINTO 62.20 52.10 57.13 42.94 70.00 84.55 61.49